OpenStudy (anonymous):

Find the derivative of m(x)=1/x+4 by using the difference quotient. m ' (x) =________ **how do you start?? thanks!!

OpenStudy (agent0smith):

Use equation editor first.

OpenStudy (agent0smith):

Or use parentheses.

OpenStudy (anonymous):

Do you know what the quotient rule?

OpenStudy (anonymous):

\[m(x)=1\div x+4\]

OpenStudy (anonymous):

is that better? and knx^(n-1) ?

OpenStudy (agent0smith):

^that doesn't help lol :P i mean use the fraction thingy

OpenStudy (anonymous):

ohh hahah lemme try to find it :P

OpenStudy (zzr0ck3r):

1/(x+4) or (1/x)+4

OpenStudy (anonymous):

\[\frac{ 1 }{ x+4 }\]

OpenStudy (zzr0ck3r):

use parentheses

OpenStudy (anonymous):

\[m(x)= 1/(x+4)\] is that better?

OpenStudy (zzr0ck3r):

yes:)

OpenStudy (anonymous):

wait how do you put the fraction thingy? :O

OpenStudy (anonymous):

in the equation tool? and what happens from here?

OpenStudy (agent0smith):

Press buttons and find out :D there's a a/b button

OpenStudy (zzr0ck3r):

you want \(\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h}=\lim_{h\rightarrow0}\frac{\frac{1}{(x+h)+4}-\frac{1}{x+4}}{h}\)

OpenStudy (zzr0ck3r):

can you simplify that?

OpenStudy (anonymous):

\[m(x)=\frac{1}{ x+4 }\]

OpenStudy (anonymous):

oh yay!

OpenStudy (anonymous):

okay um would i be multiplying by 1/h ?

OpenStudy (zzr0ck3r):

simplify the top first

OpenStudy (zzr0ck3r):

then multiply by 1/h, then some h's will cancel, then let the remaining h's be 0, then you have the answer

OpenStudy (anonymous):

ermm \[(\frac{ 1 }{ h } - 1) /\] ??

OpenStudy (anonymous):

* / h ?

OpenStudy (anonymous):

did i simplify the top correctly?

OpenStudy (dumbcow):

just use power rule with chain rule \[m = \frac{1}{x+1} = (x+1)^{-1}\] \[m' = -(1)(x+1)^{-2} = -\frac{1}{(x+1)^2}\]

OpenStudy (anonymous):

okay, so if my equation is m(x) = (1) / (x+4) it becomes... =(x+4)^-1 m ' (x) = -(1)(x+4)^-2 = - (1) / (x+4)^2 ??

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

and that's the final answer? :o

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

awesome! thank you!

OpenStudy (dumbcow):

however if for example there is an "x" on top \[\frac{x}{x+4}\] then you must use quotient rule

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

thank you!!!

OpenStudy (dumbcow):

yw